OpenStudy (anonymous):

Given f(x) = x − 7 and g(x) = x2 . Find g(f(4)). g(f(4)) =?????

OpenStudy (anonymous):

I have and I've already passed algebra 2, im doing this for a friend, and it's still not clicking in my head :\

OpenStudy (anonymous):

Did you mean \( \huge g(x) = x^2 \) ???

OpenStudy (anonymous):

Let me know if you mean \( \huge g(x) = x^2\) or \(\huge g(x) = x2 \)